OpenStudy (anonymous):

helpp!! what is the equation of the line that has a slope of 4/5 and y-intercept of -1?

OpenStudy (anonymous):

well you know this equation: y = mx + d that m is slope and d is y-intercept?

OpenStudy (anonymous):

it's equation of a line

OpenStudy (anonymous):

dont understand

OpenStudy (anonymous):

in the equation y=mx+b the m stands for the slope and the b stands for the y-int, just plug the values in to get your answer

OpenStudy (anonymous):

when we want to show a line we need a equation but in this equation we have slope and y-intercept...and we should show them in the equation so we write y = mx + d which m is SLOPE and d is Y-INTERCEPT

OpenStudy (anonymous):

4/5+1

OpenStudy (anonymous):

4/5x+1

OpenStudy (anonymous):

dont forget ur negatives!

OpenStudy (anonymous):

and y= at the front

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=4/5+1

OpenStudy (anonymous):

exactly ;)